# Can Newtonian Gravity Produce Quantum Entanglement?

### v3

@@ I Introduction
The unification of quantum mechanics and general relativity remains one of the most profound challenges in fundamental physics. In the absence of direct empirical evidence for quantum gravity at the Planck scale, recent focus has shifted toward low-energy laboratory tests that can witness the non-classicality of the gravitational field. Building on a Gedanken-experiment originally proposed by Feynman at the 1957 Chapel Hill conference [DeWittRickles2011], recent table-top proposals [Marletto:2017kzi, Bose:2017nin]—often referred to as the Bose-Marletto-Vedral (BMV) mechanism—aim to verify the quantum nature of gravity [^1] For several other experimental proposals to realize the GIE protocol, see [Marshman:2019sne, Bose:2022uxe, BeckeringVinckers:2023kco]. These proposals utilize the dynamics-independent quantum information scheme, known as the gravitationally induced entanglement (GIE) protocol. Based on the local operations and classical communication (LOCC) protocol, the GIE protocol asserts that if gravity can mediate entanglement between two spatially separated massive quantum bodies, the gravitational field itself must possess non-classical features. Consequently, the prevailing consensus within the quantum information framework is that a local classical mediator cannot generate entanglement.
A recent study [Aziz:2025ypo] posits that classical gravity, when coupled with mesoscopic quantum matter, can indeed generate entanglement, despite the classical nature of the mediating field. This claim stands in stark contrast to the dynamical constraints argued in [Marletto:2025asw, Marletto:2025fpm], which maintain that the detection of gravity-mediated entanglement constitutes a definitive probe of quantum gravity  [^2] See [Nandi:2024jyf] for a theretical model to explore the quantum nature of gravity.. The apparent contradiction between the explicit calculations in [Aziz:2025ypo] and the GIE no-go theorems [Marletto:2017kzi, Bose:2017nin] indicates a critical need to bridge the gap between abstract information-theoretic principles and concrete field-theoretic models. While the GIE argument is robust in its generality, its lack of attribution to specific exemplar gravity-matter models leaves it vulnerable to interpretational loopholes arising from technical subtleties in perturbative calculations.
@@
By studying the unitarily evolved final state of the two quantum bodies governed by quadrupole-quadrupole interactions in the three scenarios above, we conclude that the quantum bodies can be entangled in scenario (I) but not in scenarios (II) and (III). Therefore, our result is consistent with the GIE protocol. Moreover, in scenarios (II) and (III), we point out a possible loophole based on the perturbative calculation, which will yield an artifact entanglement production due to the neglect of the subleading terms by truncating the expansion at a particular order of the Newton constant. This could be a cause of the disagreement between [Aziz:2025ypo] and the GIE ones [Marletto:2017kzi, Bose:2017nin].
@@ II Preparation of mesoscopic quantum bodies
Following Feynman’s original proposal [DeWittRickles2011], a massive object may be prepared in a spatial superposition, which we call a "quantum body". Physically, this mass superposition can be realized as distributed N00N states with spatial separation of two different modes located respectively at $\vec{x}_{L}$ and $\vec{x}_{R}$ [Dowling:2008pbf, Kim:2025hbh]:
@@
where $N$ denotes the number of photons or atoms, depending on the experimental setup. Such states have recently been employed to study their quantum decoherence in the presence of a black hole [Danielson:2021egj, Danielson:2022tdw, Danielson:2022sga, Biggs:2024dgp].
@@ III Newtonian Gravity of Quantum Bodies
with $\hat{r}:={\vec{x}_{1}-\vec{x}_{2}\over|\vec{x}_{1}-\vec{x}_{2}|}$ and $r_{12}:=|\vec{x}_{1}-\vec{x}_{2}|$. This is the tidal interaction between two static and well-separated quadrupole moments, and can also be reproduced from the leading-order post-Newtonian approximation of Einstein gravity [Binnington:2009bb, Kol:2011vg, Biggs:2024dgp, Lin:2024bmg], with a possibly different $T_{ijkl}$ due to the choice of gauge fixing. Moreover, the subleading PN terms will involve the motion of the compact bodies [Goldberger:2007hy, Porto:2016pyg], which we will not consider in this note.
@@ III.1 Mini-superspace approach
where $\hat{Z}^{(a)}$ represents the $\mathbb{Z}_{2}$ parity operator for the respective quantum bodies. While this may not be a consistent treatment of quantum gravity, it is conceptually analogous to the mini-superspace approach in quantum cosmology [DeWitt:1967yk, Hartle:1983ai, Vilenkin:1984wp], which involves quantizing only the scale factor, not the full metric.
@@ III.2 Semi-classical gravity approach
In the semi-classical gravity framework [dewitt1975quantum, parker1968particle, Hawking:1975vcx], matter is quantized while the gravitational field remains classical. Consequently, the Newtonian potential is sourced by the expectation value of the quadrupole operator with respect to the N00N-like state, rather than the operator itself:
@@
Even though the Newton potential sourced by a quantum body is a classical quantity, this hybrid system exhibits bizarre behavior when a measurement collapses the quantum body, e.g., from $\theta={\pi\over 4}$ to $\theta=0$. This will then result in a sudden jump of the Newton potential due to the wavefunction collapse [Penrose:1996cv]. In general relativity, this implies that the spacetime metric can be singular.
@@ III.3 Stochastic gravity approach
Stochastic gravity extends the semi-classical approximation by incorporating fluctuations in the matter stress-energy tensor via the Einstein-Langevin equation [Hu:1994dka, Hu:2008rga, Martin:1998nc]:
@@
Here, the gravitational field is driven by a Gaussian stochastic source $\xi_{\mu\nu}$, arising from the noise kernel (fluctuations of the quantum matter) [Hu:1994dka]. This is the generalization of Langevin dynamics for Brownian motion, with the gravity field as the Brownian particle and the quantum matter as the thermally-agitated water molecules. At the Newtonian order, the potential in (10) acquires a random component corresponding to fluctuations in the quadrupole orientation. This is modeled by modifying the orientation parameter $\theta^{(a)}$:
@@ IV.1 Entanglement production in mini-superspace approach
Considering the specific initial configuration $\theta^{(a)}=\pi/4$ and $\phi^{(a)}=0$, denoted as $|{\rm init}\rangle=|++\rangle\equiv|+;1\rangle\otimes|+;2\rangle$, we define the orthogonal states $|-;a\rangle:=\hat{Z}^{(a)}|+;a\rangle$. The evolved state becomes
@@
where
$|+;a\rangle:=|\theta^{(a)}=\pi/4,\phi^{(a)}=0;Q^{(a)}_{ij}\rangle\;$. (21)
@@
for $a=1,~\text{and}~2$. Similarly, one can define $|+$-$\rangle$ and $|-+\rangle$. Then, the time-evolved final states becomes
@@ IV.3 Confusing GIE-violating result due to perturbative calculations
The recent paper [Aziz:2025ypo] claims that classical gravity can produce quantum entanglement, which is based on perturbative calculations in the context of the stochastic gravity approach. We now demonstrate a subtlety for reaching this GIE-violating result. For simplicity, we consider the initial state of (20) so that $\lambda_{a}$’s of (26) vanish as $\cos 2\theta^{(a)}\big|_{\theta^{(a)}=\pi/4}=0$ for semiclassical gravity approach, but not for stochastic gravity approach due to nontrivial $\delta\theta$. So, there is no evolution for the semiclassical gravity in this case. On the other hand, for the stochastic gravity, we have
@@
Now, we can explain the possible origin of claiming the GIE-violating result as follows. In the perturbative approach as performed in [Aziz:2025ypo], the resultant final state is obtained up to some order of $G_{N}$. This corresponds to expanding the right side of (28) in order of $\lambda_{a}$ as $\lambda_{a}\sim{\cal O}(G_{N})$. Suppose we are interested in the final state up to ${\cal O}(G_{N})$, then we have
@@ VI Acknowledgements
SM is grateful to Ignacio Araya and Giorgos Anastasiou for their hospitality at Universidad Andrés Bello and Universidad Adolfo Ibañez, respectively. The work of FLL is supported by Taiwan’s National Science and Technology Council (NSTC) through Grant No. 112-2112-M-003-006-MY3. The work of SM is supported by ANID FONDECYT Postdoctorado grant number 3240055.

### v4

@@ I Introduction
The unification of quantum mechanics and general relativity remains one of the most profound challenges in fundamental physics. In the absence of direct empirical evidence for quantum gravity at the Planck scale, recent focus has shifted toward low-energy laboratory tests that can witness the non-classicality of the gravitational field. Building on a Gedanken-experiment originally proposed by Feynman at the 1957 Chapel Hill conference [1], recent table-top proposals [2, 3]—often referred to as the Bose-Marletto-Vedral (BMV) mechanism—aim to verify the quantum nature of gravity [^1] For several other experimental proposals to realize the GIE protocol, see [4, 5, 6]. These proposals utilize the dynamics-independent quantum information scheme, known as the gravitationally induced entanglement (GIE) protocol. Based on the local operations and classical communication (LOCC) protocol, the GIE protocol asserts that if gravity can mediate entanglement between two spatially separated massive quantum bodies, the gravitational field itself must possess non-classical features. Consequently, the prevailing consensus within the quantum information framework is that a local classical mediator cannot generate entanglement.
A recent study [7] posits that classical gravity, when coupled with mesoscopic quantum matter, can indeed generate entanglement, despite the classical nature of the mediating field. This claim stands in stark contrast to the dynamical constraints argued in [8, 9], which maintain that the detection of gravity-mediated entanglement constitutes a definitive probe of quantum gravity  [^2] See [10] for a theoretical model to explore the quantum nature of gravity.. The apparent contradiction between the explicit calculations in [7] and the GIE no-go theorems [2, 3] indicates a critical need to bridge the gap between abstract information-theoretic principles and concrete field-theoretic models. While the GIE argument is robust in its generality, its lack of attribution to specific exemplar gravity-matter models leaves it vulnerable to interpretational loopholes arising from technical subtleties in perturbative calculations.
@@
By studying the unitarily evolved final state of the two quantum bodies governed by quadrupole-quadrupole interactions in the three scenarios above, we conclude that the quantum bodies can be entangled in scenario (I) but not in scenarios (II) and (III). Therefore, our result is consistent with the GIE protocol. Moreover, in scenarios (II) and (III), we point out a possible loophole based on the perturbative calculation, which will yield an artifact entanglement production due to the neglect of the subleading terms by truncating the expansion at a particular order of the Newton constant. This could be a cause of the disagreement between [7] and the GIE ones [2, 3].
@@ II Preparation of mesoscopic quantum bodies
Following Feynman’s original proposal [1], a massive object may be prepared in a spatial superposition, which we call a "quantum body". Physically, this mass superposition can be realized as distributed N00N states with spatial separation of two different modes located respectively at $\vec{x}_{L}$ and $\vec{x}_{R}$ [11, 12]:
@@
where $N$ denotes the number of photons or atoms, depending on the experimental setup. Such states have recently been employed to study their quantum decoherence in the presence of a black hole [13, 14, 15, 16].
@@ III Newtonian Gravity of Quantum Bodies
with $\hat{r}:={\vec{x}_{1}-\vec{x}_{2}\over|\vec{x}_{1}-\vec{x}_{2}|}$ and $r_{12}:=|\vec{x}_{1}-\vec{x}_{2}|$. This is the tidal interaction between two static and well-separated quadrupole moments, and can also be reproduced from the leading-order post-Newtonian approximation of Einstein gravity [17, 18, 16, 19], with a possibly different $T_{ijkl}$ due to the choice of gauge fixing. Moreover, the subleading PN terms will involve the motion of the compact bodies [20, 21], which we will not consider in this note.
@@ III.1 Mini-superspace approach
where $\hat{Z}^{(a)}$ represents the $\mathbb{Z}_{2}$ parity operator for the respective quantum bodies. While this may not be a consistent treatment of quantum gravity, it is conceptually analogous to the mini-superspace approach in quantum cosmology [22, 23, 24], which involves quantizing only the scale factor, not the full metric.
@@ III.2 Semi-classical gravity approach
In the semi-classical gravity framework [25, 26, 27], matter is quantized while the gravitational field remains classical. Consequently, the Newtonian potential is sourced by the expectation value of the quadrupole operator with respect to the N00N-like state, rather than the operator itself:
@@
Even though the Newton potential sourced by a quantum body is a classical quantity, this hybrid system exhibits bizarre behavior when a measurement collapses the quantum body, e.g., from $\theta={\pi\over 4}$ to $\theta=0$. This will then result in a sudden jump of the Newton potential due to the wavefunction collapse [28]. In general relativity, this implies that the spacetime metric can be singular.
@@ III.3 Stochastic gravity approach
Stochastic gravity extends the semi-classical approximation by incorporating fluctuations in the matter stress-energy tensor via the Einstein-Langevin equation [29, 30, 31]:
@@
Here, the gravitational field is driven by a Gaussian stochastic source $\xi_{\mu\nu}$, arising from the noise kernel (fluctuations of the quantum matter) [29]. This is the generalization of Langevin dynamics for Brownian motion, with the gravity field as the Brownian particle and the quantum matter as the thermally-agitated water molecules. At the Newtonian order, the potential in (10) acquires a random component corresponding to fluctuations in the quadrupole orientation. This is modeled by modifying the orientation parameter $\theta^{(a)}$:
@@ IV.1 Entanglement production in mini-superspace approach
For simplicity, we consider a specific initial state with $\theta^{(1)}=\theta^{(2)}={\pi\over 4}$ and $\phi^{(1)}=\phi^{(2)}=0$, i.e.,
@@
where from (2) we have defined
$|+;a\rangle:=|\theta^{(a)}=\pi/4,\phi^{(a)}=0;Q^{(a)}_{ij}\rangle$ (21)
with $a=1,2$ labeling the quantum bodies. To avoid notational clumsiness, here and below we use $|\pm;a\rangle$ to denote the eigenstates of Pauli $\hat{X}^{(a)}$ operators, and one should not confuse them with the $|\pm;Q^{(a)}_{ij}\rangle$ used in (2) for the eigenstates of $\hat{Z}^{(a)}$.
@@
Similarly, one can define $|+$-$\rangle$ and $|-+\rangle$. Then, the time-evolved final states become
@@ IV.3 Confusing GIE-violating result due to perturbative calculations
The recent paper [7] claims that classical gravity can produce quantum entanglement, which is based on perturbative calculations in the context of the stochastic gravity approach. We now demonstrate a subtlety for reaching this GIE-violating result. For simplicity, we consider the initial state of (20) so that $\lambda_{a}$’s of (26) vanish as $\cos 2\theta^{(a)}\big|_{\theta^{(a)}=\pi/4}=0$ for semiclassical gravity approach, but not for stochastic gravity approach due to nontrivial $\delta\theta$. So, there is no evolution for the semiclassical gravity in this case. On the other hand, for the stochastic gravity, we have
@@
Now, we can explain the possible origin of claiming the GIE-violating result as follows. In the perturbative approach as performed in [7], the resultant final state is obtained up to some order of $G_{N}$. This corresponds to expanding the right side of (28) in order of $\lambda_{a}$ as $\lambda_{a}\sim{\cal O}(G_{N})$. Suppose we are interested in the final state up to ${\cal O}(G_{N})$, then we have
@@ VI Acknowledgements
SM is grateful to Ignacio Araya and Giorgos Anastasiou for their hospitality at Universidad Andrés Bello and Universidad Adolfo Ibañez, respectively. The work of FLL is supported by Taiwan’s National Science and Technology Council (NSTC) through Grant No. 112-2112-M-003-006-MY3. The work of SM is supported by the ANID FONDECYT Postdoctorado grant number 3240055.